A collection of loonies, toonies, and quarters has a value of $25.50. There is one more looney than there are toonies, and there are twice as many quarters as toonies. How many of each kind of coin are in the collection?
Can you please help me?

Answer by stanbon(75887)

A collection of loonies, toonies, and quarters has a value of $25.50. There is one more looney than there are toonies, and there are twice as many quarters as toonies. How many of each kind of coin are in the collection?
------------------------------
Quantity Equation: L = T + 1
Quantity Equation: Q = 2T
Value Equation: 100L + 200T + 25Q = 2550 cents
----------------------------------------------
Substitute to solve for "T":
100(T+1) + 200T + 25(2T) = 2550
100T + 100 + 200T + 50T = 2550
350T = 2450
T = 7 (# of toonies)
Since L = T+1, L = 8 (# of loonies)
Since Q = 2L, A = 16 (# of quarters)
=========================================
Cheers,
Stan H.



===================================
Cheers,
Stan H.